solve the equation for x belongs to R x/x-1 is equal to x + 1/x-2


1
Expert's answer
2021-12-20T18:45:38-0500

"\frac{x}{x-1}=\frac{x+1}{x-2}"


"x^2-2x=x^2-1"

"2x=1"


"x=1\/2"